# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 6      $t\leftarrow t+1$
Output: Coefficients $[(\lambda_{0},\widetilde{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\widetilde{\bm{\beta}}(\lambda_{1})),\dots]$
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ and $\mathbf{y}\in\mathbb{R}^{n}$. Assume access to classical-samplable structures of size $O(n)$. Let $\delta\in(0,1)$, $\epsilon\in(0,2)$, and $T\in\mathbb{N}$. Let $\alpha,\gamma\in(0,1]$ be such that
[…]
@@
$O\left(\frac{\gamma^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}\|_{2}^{2}}{\alpha^{2}\epsilon^{2}}|\mathcal{I}|\log(Td/\delta)\operatorname{poly}\log{n}+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
We start by noticing that the joining time computation step outputs $\widetilde{i}_{t+1}^{\rm join}=\operatorname{argmax}_{i\in\mathcal{I}}\{r_{i}\}$ and $\widetilde{\lambda}_{t+1}^{\rm join}=(1-\epsilon/2)^{-1}\max_{i\in\mathcal{I}}\{r_{i}\}$, where $|r_{i}-\Lambda_{i}^{\rm join}(\mathcal{A})|\leq\epsilon\lambda_{t+1}^{\rm join}/2$. Since $\max_{i\in\mathcal{I}}\{r_{i}\}\geq r_{i_{t+1}^{\rm join}}\geq\Lambda^{\rm join}_{{i}_{t+1}^{\rm join}}(\mathcal{A})-\epsilon\lambda_{t+1}^{\rm join}/2=(1-\epsilon/2)\lambda_{t+1}^{\rm join}$, this means that the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is taken from the set
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq(1-\epsilon/2)\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\}.$
Theorem 25 thus guarantees the correctness of the algorithm. Regarding the time complexity, the most expensive steps are computing $\mathbf{X}_{\mathcal{A}}^{+}$ in time $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$, inputting $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into classical-samplable memories in time $O(n\log{n})$, and computing the $|\mathcal{I}|$ values $\{r_{i}\}_{i\in\mathcal{I}}$. By following the exact same steps as in the proof of Theorem 26, approximating any quantity $\Lambda_{i}^{\rm join}(\mathcal{A})$ within precision $\epsilon\lambda_{t+1}^{\rm join}/2$ and failure probability at most $\delta/(T|\mathcal{I}|)$ requires time
$O\left(\frac{\gamma^{-2}+n\|\mathbf{X}\|^{2}_{\max}\|\mathbf{X}^{+}\|^{2}_{2}}{\alpha^{2}\epsilon^{2}}\log(Td/\delta)\operatorname{poly}\log{n}\right)$
@@
4.4 Standard Gaussian random design matrix
The complexity of the approximate LARS algorithms depends on the quantities
$\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2},\qquad\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1},\quad\text{and}\quad\min_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{\infty}}{\|\mathbf{X}\|_{\max}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}.$
In this section, we shall bound these quantities for the specific case when the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ is the standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries. According to Fact 19, in this situation the Lasso solution is unique, thus $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and $\mathbf{X}_{\mathcal{A}}^{+}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top}$. As we shall prove below, with high probability, $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$ is a constant and the mutual incoherence is at least $1/2$. The mutual overlap between $\mathbf{y}$ and $\mathbf{X}$, on the other hand, is $\Omega(1/\sqrt{n\log{d}})$ with high probability, which preserves the overall $\widetilde{O}(\sqrt{n|\mathcal{I}|})$ complexity of the approximate quantum LARS algorithm.
Let us start our analysis with the quantity $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$. The next result states that it is basically a constant with high probability in the high-dimensional case.
@@ 4.4 Standard Gaussian random design matrix
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ with $d\geq 2n$ be a standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries. Then, for any $\delta\in(0,1)$, $\mathbb{P}[\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}\leq e^{3/\delta}]\geq 1-\delta$.
@@
In order to bound the quantity $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$, first note that it is upper bounded by the condition number $\kappa(\mathbf{X})\triangleq\|\mathbf{X}\|_{2}\|\mathbf{X}^{+}\|_{2}$ of $\mathbf{X}$. The condition number of standard Gaussian random matrices was studied by Chen and Dongarra [CD05], who proved that $\mathbb{E}[\ln\kappa(\mathbf{X})]\leq 2.258+\ln\frac{d}{|d-n|+1}$ (see also [ES05]). Thus, in the case when $d\geq 2n$, $\mathbb{E}[\ln\kappa(\mathbf{X})]\leq 2.258+\ln{2}\leq 3$. Hence, by Markov’s inequality, $\mathbb{P}[\kappa(\mathbf{X})\leq e^{3/\delta}]\geq 1-\delta$ for any $\delta\in(0,1)$. ∎
We now turn our attention to
$\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}=\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}=\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\|\mathbf{X}_{\mathcal{A}^{c}}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{\infty}.$
We prove in the next lemma that, if $T=O(\sqrt{n/\log{d}})$, then the mutual incoherence is a constant bounded away from 00 with high probability. Our result applies to the more general case when the rows of $\mathbf{X}$ are sampled i.i.d. according to a $\mathcal{N}(0,\mathbf{\Sigma})$ distribution, where $\mathbf{\Sigma}\in\mathbb{R}^{d\times d}$. A similar result can be found in [Wai19, Exercise 7.19].
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a random matrix with rows sampled i.i.d. according to a $\mathcal{N}(0,\mathbf{\Sigma})$ distribution, where $\mathbf{\Sigma}\in\mathbb{R}^{d\times d}$. Let $\delta\in(0,1)$ and $\gamma_{\min}(\mathbf{\Sigma})\in(0,1]$ be the minimum eigenvalue of $\mathbf{\Sigma}$. Suppose that the diagonal entries of $\mathbf{\Sigma}$ are at most $1$ and that there is $\bar{\alpha}\in(0,1]$ such that
$\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\max_{j\in\mathcal{A}^{c}}\|\mathbf{\Sigma}_{j\mathcal{A}}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})^{-1}\|_{1}\leq 1-\bar{\alpha},$
for some $T\in\mathbb{N}$. If $T=O\big{(}\sqrt{n\bar{\alpha}^{2}\gamma_{\min}(\mathbf{\Sigma})/\log(d/\delta)}\big{)}$, then
$\mathbb{P}\left[\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\|\mathbf{X}_{\mathcal{A}^{c}}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{\infty}\leq 1-\frac{\bar{\alpha}}{2}\right]\geq 1-\delta.$
@@
For fixed $\mathcal{A}$ and any $\mathbf{u}\in\mathbb{R}^{|\mathcal{A}|}$ such that $\|\mathbf{u}\|_{\infty}=1$, first note that
@@
Consider the quantity $|\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|$ for fixed $j\in\mathcal{A}^{c}$. The zero-mean Gaussian vector $\mathbf{X}_{j}$ can be decomposed into a linear prediction based on $\mathbf{X}_{\mathcal{A}}$ plus a prediction error as (see [Wai09, Section V.A] and [Wai19, Exercise 11.3])
where $\mathbf{W}_{j}\in\mathbb{R}^{n}$ is a vector with i.i.d. $\mathcal{N}(0,[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj})$ entries. Here the (positive semi-definite) matrix $\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}=\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}^{c}}-\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})^{-1}\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}^{c}}$ is the conditional co-variance matrix of $(\mathbf{X}_{\mathcal{A}^{c}}|\mathbf{X}_{\mathcal{A}})$. Thus
@@
$\leq 1-\bar{\alpha}+|\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|.$
Since $\mathbf{W}_{j}$ is independent of $\mathbf{X}_{\mathcal{A}}$, $\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}$ is a sub-Gaussian random variable with parameter (Fact 6)
@@
where we used that $[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\leq 1$ since $\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}$ is positive semi-definite and the diagonal entries of $\mathbf{\Sigma}$ are at most $1$. We now relate the quantity $\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}$ with $\gamma_{\min}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})\geq\gamma_{\min}(\mathbf{\Sigma})$. According to [Wai09, Lemma 9],
@@
$\mathbb{P}\left[\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}\geq\frac{(1+t+\sqrt{|\mathcal{A}|/n})^{2}}{n\gamma_{\min}(\mathbf{\Sigma})}\right]\leq 2e^{-nt^{2}/2}\quad\text{for all}~{}t>0.$

### v2

@@ 23      $t\leftarrow t+1$
return coefficients $[(\lambda_{0},\widetilde{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\widetilde{\bm{\beta}}(\lambda_{1})),\dots]$
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ and $\mathbf{y}\in\mathbb{R}^{n}$. Assume access to classical-samplable structures of size $O(n)$. Let $\delta\in(0,1)$, $\epsilon>0$, and $T\in\mathbb{N}$. For $\mathcal{A}\subseteq[d]$ of size $|\mathcal{A}|\leq T$, assume there are $\alpha_{\mathcal{A}},\gamma_{\mathcal{A}}\in(0,1)$ such that
$\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\leq\alpha_{\mathcal{A}}\quad\text{and}\quad\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{\infty}}{\|\mathbf{X}\|_{\max}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}\geq\gamma_{\mathcal{A}}.$
@@
$O\left(\frac{\gamma_{\mathcal{A}}^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}^{2}}{(1-\alpha_{\mathcal{A}})^{2}\epsilon^{2}}|\mathcal{I}|\log(Td/\delta)\operatorname{poly}\log{n}+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
We start by noticing that the joining time computation step outputs $\widetilde{i}_{t+1}^{\rm join}=\operatorname{argmax}_{i\in\mathcal{I}}\{r_{i}\}$ and $\widetilde{\lambda}_{t+1}^{\rm join}=\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}^{-1}\max_{i\in\mathcal{I}}\{r_{i}\}$, where $|r_{i}-\Lambda_{i}^{\rm join}(\mathcal{A})|\leq\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\lambda_{t+1}^{\rm join}$. Since $\max_{i\in\mathcal{I}}\{r_{i}\}\geq r_{i_{t+1}^{\rm join}}\geq\Lambda^{\rm join}_{{i}_{t+1}^{\rm join}}(\mathcal{A})-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\lambda_{t+1}^{\rm join}=\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}\lambda_{t+1}^{\rm join}$, this means the joining variable $\widetilde{i}_{t+1}^{\rm join}$ is taken from the set
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\}.$
Theorem 28 thus guarantees the correctness of the algorithm. Regarding the time complexity, the most expensive steps are computing $\mathbf{X}_{\mathcal{A}}^{+}$ in time $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$, inputting $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into classical-samplable memories in time $O(n\log{n})$, and computing the $|\mathcal{I}|$ values $\{r_{i}\}_{i\in\mathcal{I}}$. By following the exact same steps as in the proof of Theorem 29, approximating any quantity $\Lambda_{i}^{\rm join}(\mathcal{A})$ within precision $\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\lambda_{t+1}^{\rm join}$ and failure probability at most $\frac{\delta}{T|\mathcal{I}|}$ requires time
$O\left(\frac{\gamma_{\mathcal{A}}^{-2}+n\|\mathbf{X}\|^{2}_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|^{2}_{2}}{(1-\alpha_{\mathcal{A}})^{2}\epsilon^{2}}\log(Td/\delta)\operatorname{poly}\log{n}\right)$
@@
If $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}=o(1)$ and the parameters $\alpha_{\mathcal{A}}$ and $\gamma_{\mathcal{A}}$ are bounded away from $1$ and 00, respectively, then it is possible to obtain a speedup compared to the standard LARS algorithm as we shall see in the next section. For several inputs $\mathbf{X}$ and $\mathbf{y}$, however, we expect Algorithm 4 to perform worse than Algorithm 1. We interpret this as evidence for the difficulty in properly dequantising the approximate quantum LARS Algorithm 3. For ill-behaved design matrices, approximating a function of its entries should be as costly as exactly computing it.
4.4 Gaussian random design matrix
For $\mathcal{A}\subseteq[d]$, the complexity of the approximate LARS algorithms depends on the quantities
$\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2},\qquad\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1},\qquad\text{and}\qquad\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{\infty}}{\|\mathbf{X}\|_{\max}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}.$
In this section, we shall bound these quantities for the specific case when the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ is a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble. According to Fact 22, in this situation the Lasso solution is unique almost surely, thus $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ and $\mathbf{X}_{\mathcal{A}}^{+}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top}$. As we shall prove below, with high probability, $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$ is $O(\sqrt{\log(d)/n)}$ and the mutual incoherence is at most $1/2$. The mutual overlap between $\mathbf{y}$ and $\mathbf{X}$, on the other hand, is $\Omega(\|\mathbf{y}\|_{2}/(\|\mathbf{y}\|_{1}\sqrt{\log{d}}))$ with high probability, which equals $\Omega(1/\sqrt{\log{d}})$ when $\mathbf{y}$ is sparse. These bounds yield quantum and classical complexities $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$ per iteration for Algorithm 3 and Algorithm 4, respectively, as stated in Result 4.
Let us start our analysis with the quantity $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$. The next result states that, if $\mathbf{\Sigma}$ is well behaved, which includes $\mathbf{\Sigma}=\mathbf{I}$, then $\|\mathbf{X}\|_{\max}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2}=O(\sqrt{\log(d)/n})$ with high probability.
@@ 4.4 Gaussian random design matrix
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble, $\mathbf{\Sigma}\succ\mathbf{0}$. Given $\mathcal{A}\subseteq[d]$ of size $|\mathcal{A}|\leq n$, let $\mathbf{\Sigma}_{\mathcal{A}|\mathcal{A}^{c}}=\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}}-\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}^{c}}(\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}^{c}})^{-1}\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}}\succ\mathbf{0}$ be the conditional covariance matrix of $\mathbf{X}_{\mathcal{A}}|\mathbf{X}_{\mathcal{A}^{c}}$. Then, for $\delta\in(2e^{-n/2},1)$,
[…]
In particular, if $\mathbf{\Sigma}=\mathbf{I}$ and $|\mathcal{A}|\leq n/2$, then
$\mathbb{P}\left[\|\mathbf{X}\|_{\max}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2}\leq\frac{\sqrt{\ln(4nd/\delta)}}{(1-1/\sqrt{2})\sqrt{n}-\sqrt{2\ln(2/\delta)}}\right]\geq 1-\delta.$
@@
Let us start by bounding $\|\mathbf{X}\|_{\max}$. We can write $\mathbf{X}=\mathbf{Z}\sqrt{\mathbf{\Sigma}}$ where $\mathbf{Z}$ is drawn from the $\mathbf{I}$-Gaussian ensemble. Let $\mathbf{Z}^{(i)}\in\mathbb{R}^{d\times 1}$ be the $i$-th row of $\mathbf{Z}$. Then
$\|\mathbf{X}\|_{\max}=\|\mathbf{Z}\sqrt{\mathbf{\Sigma}}\|_{\max}=\max_{i\in[n]}\|\sqrt{\mathbf{\Sigma}}^{\top}\mathbf{Z}^{(i)}\|_{\infty}\leq\max_{i\in[n]}\|\sqrt{\mathbf{\Sigma}}^{\top}\|_{\infty}\|\mathbf{Z}^{(i)}\|_{\infty}=\|\sqrt{\mathbf{\Sigma}}\|_{1}\|\mathbf{Z}\|_{\max}.$
By a Chernoff bound (Fact 6) plus a union bound, we have that $\mathbb{P}[\|\mathbf{Z}\|_{\max}\leq\sqrt{\ln(4nd/\delta)}]\geq 1-\delta/2$, which implies that $\mathbb{P}[\|\mathbf{X}\|_{\max}\leq\sqrt{\ln(4nd/\delta)}\|\sqrt{\mathbf{\Sigma}}\|_{1}]\geq 1-\delta/2$.
We now focus on $\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$. The rows of $\mathbf{X}_{\mathcal{A}}$ are i.i.d. sampled from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{\mathcal{A}|\mathcal{A}^{c}})$. According to [114, Theorem 6.1],
[…]
By taking $\epsilon=\sqrt{2\ln(2/\delta)/n}<1$, the above probability is at most $\delta/2$. The result follows by putting together the bounds on $\|\mathbf{X}\|_{\max}$ and $\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2}$.
[…]
We now move on to $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}$. The next lemma tells us that, if $|\mathcal{A}|=O(\sigma_{\min}(\mathbf{\Sigma})n/\log{d})$, where $\sigma_{\min}(\mathbf{\Sigma})$ is the minimum eigenvalue of $\mathbf{\Sigma}$, then the mutual incoherence is a constant bounded away from $1$ with high probability. A similar result can be found in [114, Exercise 7.19].
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble, where $\mathbf{\Sigma}\succ\mathbf{0}$ has minimum eigenvalue $\sigma_{\min}(\mathbf{\Sigma})\in(0,1]$ and its diagonal entries are at most $1$. Given $\mathcal{A}\subseteq[d]$, suppose that there is $\bar{\alpha}\in[0,1)$ such that
$\max_{j\in\mathcal{A}^{c}}\|\mathbf{\Sigma}_{j\mathcal{A}}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})^{-1}\|_{1}\leq\bar{\alpha}.$
Given $\delta\in(0,1)$, if $|\mathcal{A}|\leq(1-\bar{\alpha})^{2}\sigma_{\min}(\mathbf{\Sigma)}n/72\ln(4d/\delta)$, then
$\mathbb{P}\left[\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\leq\frac{1}{2}+\frac{\bar{\alpha}}{2}\right]\geq 1-\delta.$
@@
First notice that
$\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}=\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}=\|\mathbf{X}_{\mathcal{A}^{c}}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{\infty}.$
For any $\mathbf{u}\in\mathbb{R}^{|\mathcal{A}|}$ such that $\|\mathbf{u}\|_{\infty}=1$, note that
@@
Consider the quantity $|\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|$ for fixed $j\in\mathcal{A}^{c}$. The zero-mean Gaussian vector $\mathbf{X}_{j}$ can be decomposed into a linear prediction based on $\mathbf{X}_{\mathcal{A}}$ plus a prediction error as (cf. [113, Section V.A] and [114, Equation (11.26) & Exercise 11.3])
$\mathbf{X}_{j}^{\top}=\mathbf{\Sigma}_{j\mathcal{A}}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top}+\mathbf{W}_{j}^{\top},$ (15)
where $\mathbf{W}_{j}\in\mathbb{R}^{n}$ is a vector with i.i.d. $\mathcal{N}(0,[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj})$ entries. Here the (positive-definite) matrix $\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}=\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}^{c}}-\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})^{-1}\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}^{c}}$ is the conditional covariance matrix of $\mathbf{X}_{\mathcal{A}^{c}}|\mathbf{X}_{\mathcal{A}}$. Thus
@@
$\leq\bar{\alpha}+|\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|.$
Since $\mathbf{W}_{j}$ is independent of $\mathbf{X}_{\mathcal{A}}$, $\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}$ is a sub-Gaussian random variable with parameter (Fact 7)
@@
where we used that $[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\leq 1$ since $\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}\preceq\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}^{c}}$ and the diagonal entries of $\mathbf{\Sigma}$ are at most $1$. We now relate $\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}$ with $\sigma_{\min}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})\geq\sigma_{\min}(\mathbf{\Sigma})$. According to [113, Lemma 9],
@@
$\mathbb{P}\left[\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}\geq\frac{(1+t+\sqrt{|\mathcal{A}|/n})^{2}}{n\sigma_{\min}(\mathbf{\Sigma})}\right]\leq 2e^{-nt^{2}/2}\quad\text{for all}\leavevmode\nobreak\ t>0.$